OpenStudy (anonymous):

What is the best order-of-magnitude estimate for 58,100? 10^3 10^4 10^5 10^6

OpenStudy (cherokeegutierrez):

the second one

OpenStudy (anonymous):

10^3=1,000-9,999 which is 4 digits 10^4=10,000-99,999 which is 5 digits 10^5=100,000-999,999 which is 6 digits 10^6=1,000,000-9,999,999 which is 7 digits

OpenStudy (koikkara):

\(10^4 = 10,000 \approx 58,100\) while, \(10^5\) is the closer one to 58,100. @pmnodine \(\color{blue}{\huge\tt{Nice~}}\) \(\huge\mathcal{to~meet~}\) \(\color{red}{\huge\tt{You~!}}\)

OpenStudy (anonymous):

No its not. @Koikkara It is asking for orders of magnitude not "which number here is closest to 58,100?"

OpenStudy (anonymous):

Magnitude=digits.

OpenStudy (koikkara):

hmm... I \(don't~ know\) how to explain... \(10^4=10,000\) \(10^5=1,00,000\) but \(58,100\) comes in the order of digits of this similiar fashion \(10,000\) not \(1,00,000\) @Shalante

OpenStudy (anonymous):

\[10^{4}=10,000\] \[10^{5}=100,000\] You just said it comes in similar fashion to 10,000 which is to the 4th power. Thats what I said earlier, but you mentioned 10^5